find the domain of 3x/x^2+1
Answer by malakumar_kos@yahoo.com(315)

find the domain of 3x/x^2+1

f(x) = 3x/x^2+1 the values of x are called range and the corresponding values of f(x) are called domain.
let us take the values of x as {0,1,2,3,4,5}
if x = 0 then f(x) = 3(0)/0^2+1 = 0/1 = 0
if x = 1 then f(x) = 3(1)/1^2+1 = 3/1+1= 3/2
if x= 2 then f(x) = 3(2)/2^2+1 = 6/4+1 = 6/5
if x= 3 then f(x) = 3(3)/3^2+1 = 9/9+1 = 9/10

if x= 4 then f(x) = 3(4)/4^2+1 = 12/16+1 = 12/17
if x= 5 then f(x) = 3(5)/5^2+1 = 15/25+1 = 15/26
the domain is {0,3/2,6/5,9/10,12/17,15/26}